# Multi-Agent Taskforce Collaboration: Self-Correction of Compounding Errors in Long-Form Literature Review Generation

### v1

@@ Introduction
We first evaluate our framework on the AutoSurvey (Wang et al. 2024b) benchmark, following the protocols established by Liang et al. (2025); Wang et al. (2024b). We use the same 20 topics from diverse subfields of LLM research to generate survey articles for comparison. We compare our approach with three baselines: Naive RAG-based LLMs using retrieval-augmented generation, AutoSurvey (Wang et al. 2024b), and SurveyX (Liang et al. 2025). We use the official implementations of AutoSurvey [^1] https://github.com/AutoSurveys/AutoSurvey and SurveyX [^2] https://github.com/IAAR-Shanghai/SurveyX to conduct experiments. Due to the lack of an online version of SurveyX, we conduct offline generation.
@@
We further evaluate our framework on the SurveyEval benchmark (Wang et al. 2025b), using the same protocols. SurveyEval is the first benchmark in computer science that pairs surveys with complete reference papers, comprising 384 arXiv cs.CL surveys citing over 26,000 references. Twenty topics were selected for testing based on reference completeness and reference list diversity. Experimental results are shown in Fig. 3, with Coverage, Structure, and Relevance scores normalized to a 100-point scale.
Methods Citation Quality Content Quality Rec. $\uparrow$ Pre. $\uparrow$ Cov. $\uparrow$ Str. $\uparrow$ Rel. $\uparrow$ Avg. $\uparrow$ w/o exploration 94.38 88.56 4.83 4.83 5.00 4.89 w/o exploitation 97.86 79.02 4.97 4.93 4.93 4.94 w/o experience 97.78 80.82 4.79 4.89 4.97 4.88 Proposed 98.17 89.28 4.97 4.95 5.00 4.97
@@
Methods Citation Quality Content Quality Recall $\uparrow$ Precision $\uparrow$ Coverage $\uparrow$ Structure $\uparrow$ Relevance $\uparrow$ Avg. $\uparrow$ Human 93.07 87.76 5.00 4.97 5.00 4.99 Naive RAG (2024) 64.57 61.89 4.29 3.58 4.67 4.18 AutoSurvey (2024) 70.03 71.66 4.68 4.67 4.87 4.74 SurveyX (2025) 75.51 77.90 4.71 4.84 4.93 4.83 Proposed 86.63 81.98 4.85 4.90 5.00 4.92
@@
Method Iteration Citation Quality Content Quality Rec. $\uparrow$ Pre. $\uparrow$ Cov. $\uparrow$ Str. $\uparrow$ Rel. $\uparrow$ Avg. $\uparrow$ Exper. 1 83.64 74.37 4.61 4.79 4.94 4.78 2 84.99 78.19 4.77 4.83 4.98 4.86 3 85.11 80.13 4.82 4.86 4.98 4.89 4 86.63 81.98 4.85 4.90 5.00 4.92 5 86.71 81.86 4.85 4.88 4.98 4.90 Explor. 1 82.84 81.26 4.71 4.78 5.00 4.83 2 84.79 81.67 4.83 4.82 5.00 4.88 3 85.69 81.97 4.85 4.84 5.00 4.90 4 86.63 81.98 4.85 4.90 5.00 4.92 5 86.65 81.92 4.85 4.87 5.00 4.91 Exploi. 1 86.31 72.68 4.81 4.87 4.91 4.86 2 86.00 77.34 4.83 4.90 4.94 4.89 3 86.75 79.10 4.84 4.89 4.98 4.90 4 86.63 81.98 4.85 4.90 5.00 4.92 5 86.47 80.27 4.84 4.89 5.00 4.91
@@
We further conduct an ablation study to evaluate the effectiveness of each component in our proposed framework. Experiments are performed on the same benchmark dataset as before, with all experimental settings unchanged and the survey length set to 8k tokens. We compare the full model with three variants. We use a single round of retrieval and organization for the variant without the exploration taskforce to generate the overall outline. We draft the manuscript using only one round of extraction and writing for the variant without the exploitation taskforce. For the variant without the experience taskforce, each agent completes its assigned task, without revision.
@@
Table 2 presents the results. Removing the exploration taskforce causes the largest drops in recall and structure, demonstrating its importance for citation recall and logical organization. Excluding the exploitation taskforce sharply reduces precision and relevance, confirming its role in improving citation precision and content relevance. Without the experience taskforce, all metrics decrease, especially coverage, highlighting its key role in ensuring comprehensive coverage through collaborative revision.
A New Large-Scale Benchmark
@@
To further validate the effectiveness of our proposed framework, we construct a new large-scale benchmark dataset. This dataset consists of 195 topics from various subfields of computer science, nearly 10 times larger than previous benchmarks (Wang et al. 2024b; Liang et al. 2025). To ensure topic quality, we collect peer-reviewed survey topics from top computer science conferences, rather than from preprint sources such as arXiv. Survey papers accepted only as abstracts are excluded. To prevent data leakage from LLM pretraining data, we include only survey papers published in 2023, 2024, and 2025. We employ PhD students in computer science to verify whether a paper qualifies as a survey and to collect the final set of 195 survey papers: 9 from AAAI, 34 from ACL, 46 from EMNLP, 3 from ICLR, 2 from ICML, 77 from IJCAI, and 24 from NAACL. Of these, 68 were published in 2023, 105 in 2024, and 22 in 2025.
@@
We conduct experiments using the same settings as above, generating 64k-token literature reviews for evaluation. As shown in Table 3, compared with results on the existing benchmark in Table 1, performance on the new large-scale benchmark drops significantly, particularly in recall and coverage. This may be due to the greater number and broader range of topics, which leads to some relevant literature not being retrieved, resulting in lower citation recall and coverage. Precision also drops slightly, likely because the models lack knowledge of the most recent two years. Therefore, compared to existing benchmarks, the new large-scale benchmark is more challenging and leads to more errors in generation. Despite this, our proposed method achieves over 80% in citation scores and an average content quality score of 4.92.
Sensitivity Analysis
@@
To gain deeper insight into our framework, we conduct a sensitivity analysis to observe how self-improving iterations affect performance. We evaluate the experience, exploration, and exploitation taskforces across different numbers of iterations, employing a controlled variable approach: when varying one hyperparameter, all others are held constant as previously described.
@@
Conclusion
In this paper, we tackle the challenge of compounding errors in automatic literature review generation by introducing Multi-agent Taskforce Collaboration (MATC). We propose three taskforces—exploration, exploitation, and experience—to mitigate both inter- and intra-agent compounding errors. Extensive experiments demonstrate that MATC outperforms existing methods on established benchmarks as well as on our contributed large-scale benchmark. Future work will collect fine-grained human annotations for better evaluation of the proposed method. We also plan to extend MATC to mitigate errors in automatic experimentation and automatic review, further benefiting AI for science (Zhang et al. 2025a; Chen et al. 2025).

### v2

@@ 4.2 Comparision Experiments
We first evaluate our framework on the AutoSurvey Wang et al. (2024b) benchmark, following the protocols established by Liang et al. (2025); Wang et al. (2024b). We use the same 20 topics from diverse subfields of LLM research to generate survey articles for comparison. We compare our approach with three baselines: Naive RAG-based LLMs using retrieval-augmented generation, AutoSurvey Wang et al. (2024b), and SurveyX Liang et al. (2025). We follow the official report of AutoSurvey [^1] https://github.com/AutoSurveys/AutoSurvey and conduct experiments with the official implementation of the offline version’s SurveyX [^2] https://github.com/IAAR-Shanghai/SurveyX.
@@
We further evaluate our framework on the SurveyEval benchmark Wang et al. (2025b), using the same protocols. SurveyEval is the first benchmark in computer science that pairs surveys with complete reference papers, comprising 384 arXiv cs.CL surveys citing over 26,000 references. Twenty topics were selected for testing based on reference completeness and reference list diversity. Experimental results are shown in Fig. 3, with Coverage, Structure, and Relevance scores normalized to a 100-point scale.
[…]
@@
[…]
@@
[…]
@@ 4.3 Ablation Study
We further conduct an ablation study to evaluate the effectiveness of each component in our proposed framework. Experiments are performed on the same benchmark dataset as before, with all experimental settings unchanged and the survey length set to 8k tokens. We compare the full model with three variants. We use a single round of retrieval and organization for the variant without the exploration taskforce to generate the overall outline. We draft the manuscript using only one round of extraction and writing for the variant without the exploitation taskforce. For the variant without the feedback taskforce, each agent completes its assigned task, without revision.
@@
Table 2 presents the results. Removing the exploration taskforce causes the largest drops in recall and structure, demonstrating its importance for citation recall and logical organization. Excluding the exploitation taskforce sharply reduces precision and relevance, confirming its role in improving citation precision and content relevance. Without the feedback taskforce, all metrics decrease, especially coverage, highlighting its key role in ensuring comprehensive coverage through collaborative revision.
4.4 A New Large-Scale Benchmark
@@
To further validate the effectiveness of our proposed framework, we construct a new large-scale benchmark dataset. This dataset consists of 195 topics from various subfields of computer science, nearly 10 times larger than previous benchmarks Wang et al. (2024b); Liang et al. (2025). To ensure topic quality, we collect peer-reviewed survey topics from top computer science conferences, rather than from preprint sources such as arXiv. Survey papers accepted only as abstracts are excluded. To prevent data leakage from LLM pretraining data, we include only survey papers published in 2023, 2024, and 2025. We employ PhD students in computer science to verify whether a paper qualifies as a survey and to collect the final set of 195 survey papers: 9 from AAAI, 34 from ACL, 46 from EMNLP, 3 from ICLR, 2 from ICML, 77 from IJCAI, and 24 from NAACL. Of these, 68 were published in 2023, 105 in 2024, and 22 in 2025.
@@
We conduct experiments using the same settings as above, generating 64k-token literature reviews for evaluation. As shown in Table 3, compared with results on the existing benchmark in Table 1, performance on the new large-scale benchmark drops significantly, particularly in recall and coverage. This may be due to the greater number and broader range of topics, which leads to some relevant literature not being retrieved, resulting in lower citation recall and coverage. Despite this, our proposed method achieves over 80% in citation scores and an average content quality score of 4.92.
To verify the consistency between LLM evaluation methods and human evaluation, we follow Wang et al. (2024b) to conduct a meta-evaluation. Human experts rank the generated surveys, and we compare these rankings with those generated by LLMs using Spearman’s rank correlation coefficient. The LLM evaluation achieves a correlation of 0.5117, suggesting LLMs align well with human preferences.
4.5 Sensitivity Analysis
@@
To gain deeper insight into our framework, we conduct a sensitivity analysis to observe how $d_{\max}$, $t_{\max}$, and $x_{\max}$ affect performance. We evaluate the experience, exploration, and exploitation taskforces across different numbers of iterations, employing a controlled variable approach: when varying one hyperparameter, all others are held constant as previously described.
@@ 4.6 Disscussions
5 Conclusion
In this paper, we identified and addressed the critical challenge of compounding errors in automated long-form literature review generation. The proposed Multi-Agent Taskforce Collaboration framework strategically orchestrates LLM-based agents through three specialized taskforces. Extensive experiments demonstrate that MATC achieves state-of-the-art performance across established benchmarks. The robustness and generalizability of our framework are further confirmed on our newly contributed TopSurvey benchmark, a large-scale collection of 195 peer-reviewed topics. The deployment of our online system, which has generated over 20,000 reviews, provides practical evidence of MATC’s effectiveness and efficiency in real-world scenarios.